OpenStudy (anonymous):

Explain the pythagorean theorem and how we would find the hypotenuse of a triangle with legs of 6 and 8.

OpenStudy (anonymous):

square, add, take the square root

Parth (parthkohli):

For the sake of complicating it: Define three functions: \(f:x\mapsto\sqrt{x}\). \(g:x \mapsto x^2\) \(t:(a,b)\mapsto a + b\) Then your answer is \(f(t(g(6,8)))\).

Parth (parthkohli):

Seem difficult, eh?

OpenStudy (anonymous):

a bit

Parth (parthkohli):

Basically,\[c^2 = a^2 + b^2\]Where \(c\) is the hypotenuse and \(a,b\) are legs. We know that \(6,8\) are the legs.

Parth (parthkohli):

So just a matter of solving \(c^2 = 6^2 + 8^2\) for \(c\).